a 5 foot border is added to all four sides of a rectangular playing field of 40 feet by 30 feet.What is the area in square feet of the added border?
Answer by checkley77(12844)

40*30=1,200 ft^2 is the area of the playing field.
(40+10)(30+10)=50*40=2,000 ft^2 is the total area of the field & the border.
2,000-1,200=800ft^2 is the area of the border.